# Spurious Correlations in High Dimensional Regression: The Roles of Regularization, Simplicity Bias and Over-Parameterization

### v1

@@ B.1 Proofs on $S^{\Sigma}_{x}$
$=x^{\top}\left(\Sigma_{xx}^{-1}+\Sigma_{xx}^{-1}\Sigma_{xy}{S^{\Sigma}_{x}}^{-1}\Sigma_{yx}\Sigma_{xx}^{-1}\right)x+y^{\top}{S^{\Sigma}_{x}}^{-1}y\,+$
$\qquad+x^{\top}\left(-\Sigma_{xx}^{-1}\Sigma_{xy}{S^{\Sigma}_{x}}^{-1}\right)y+y^{\top}\left(-{S^{\Sigma}_{x}}^{-1}\Sigma_{yx}\Sigma_{xx}^{-1}\right)x$
@@ B.2 Remarks on Assumption 1
Let Assumptions 4 and 6 hold, and let $n=\Theta(d)$. Let $\hat{\theta}_{\textup{LR}}(\lambda)$ be defined as in (4.4), and let $\mathcal{C}(\hat{\theta}_{\textup{LR}}(\lambda))$ be the amount of spurious correlations learned by the model $f_{\textup{LR}}\left(\hat{\theta}_{\textup{LR}}(\lambda)\right)$ as defined in (3.3). Then, for any $\lambda>0$, we have that, for every $t\in(0,1/2)$,
@@
Let Assumptions 4 and 6 hold, and let $n=\Theta(d)$. Let $\hat{\theta}_{\textup{LR}}(\lambda)$ be defined as in (4.4), and let $\mathcal{L}(\hat{\theta}_{\textup{LR}}(\lambda))$ be the in-distribution test loss of the model $f_{\textup{LR}}(\hat{\theta}_{\textup{LR}}(\lambda))$ as defined in (4.1). Then, , for any $\lambda>0$, we have that, for every $t\in(0,1/2)$,
@@ Appendix C Proofs for Random Features
$\begin{split}\lambda_{\rm min}\left(\Phi\Phi^{\top}\right)&\geq p\,\lambda_{\rm min}\left(\mu_{1}^{2}\frac{ZZ^{\top}}{2d}+{\tilde{\mu}}^{2}I\right)-\left\lVert\mathbb{E}_{V}\left[\Phi\Phi^{\top}\right]-p\left(\mu_{1}^{2}\frac{ZZ^{\top}}{2d}+{\tilde{\mu}}^{2}I\right)\right\rVert_{\textup{op}}\\ &\qquad-\left\lVert\Phi\Phi^{\top}-\mathbb{E}_{V}\left[\Phi\Phi^{\top}\right]\right\rVert_{\textup{op}}\\ &\geq p{\tilde{\mu}}^{2}-\mathcal{O}\left(\frac{p\log^{3}d}{\sqrt{d}}\right)-\mathcal{O}\left(\sqrt{pd}\right)=\Omega(p),\end{split}$ (C.30)
@@
Let $z\in\mathbb{R}^{2d}$ be sampled from a distribution respecting Assumption 4, not necessarily with the same covariance as $\mathcal{P}_{XY}$, independent from everything else, and let $f_{\textup{RF}}(z,\hat{\theta}_{\textup{RF}}(\lambda))$ be the RF model defined in (6.1) with $\hat{\theta}_{\textup{RF}}(\lambda)$ defined in (6.3), with $\lambda\geq 0$. Then, we have that
$\left|f_{\textup{RF}}(z,\hat{\theta}_{\textup{RF}}(\lambda))-\mu_{1}^{2}p\frac{z^{\top}Z^{\top}}{2d}\left(\Phi\Phi^{\top}+n\lambda I\right)^{-1}G\right|=\mathcal{O}\left(\frac{d^{1/4}\log d}{p^{1/4}}+\frac{\log^{3/2}d}{d^{1/8}}\right)=o(1),$ (C.58)
@@
$f_{\textup{RF}}(z,\hat{\theta}_{\textup{RF}}(\lambda))$ $=\left(\mu_{1}Vz+\tilde{\phi}\left(Vz\right)\right)^{\top}\left(\mu_{1}VZ^{\top}+\tilde{\Phi}^{\top}\right)\hat{G}$ (C.59)
@@
$\left|f_{\textup{RF}}(z,\hat{\theta}_{\textup{RF}}(\lambda))-\mu_{1}^{2}p\frac{z^{\top}Z^{\top}}{2d}\hat{G}\right|=\mathcal{O}\left(\frac{d^{1/4}\log d}{p^{1/4}}+\frac{\log^{3/2}d}{d^{1/8}}\right),$ (C.68)
@@
$=f_{\textup{LR}}(z,\hat{\theta}_{\textup{LR}}(\tilde{\lambda})),$
@@
$\left|\mu_{1}^{2}p\frac{z^{\top}Z^{\top}}{2d}\left(\Phi\Phi^{\top}+n\lambda I\right)^{-1}G-f_{\textup{RF}}(z,\hat{\theta}_{\textup{RF}}(\lambda))\right|=\mathcal{O}\left(\frac{d^{1/4}\log d}{p^{1/4}}+\frac{\log^{3/2}d}{d^{1/8}}\right),$ (C.75)
@@
Appendix D Proof of (3.4): connection with out-of-distribution loss
@@
$\mathbb{E}_{\tilde{x},y}\left[\left(f(\hat{\theta},[\tilde{x}^{\top},y])-f^{*}_{x}(\tilde{x})\right)^{2}\right]$ $=2-2\mathbb{E}_{\tilde{x},y}\left[f(\hat{\theta},[\tilde{x}^{\top},y])f^{*}_{x}(\tilde{x})\right]$ (D.1)
@@
$S=\left(\begin{array}[]{@{}c c c@{}}1&\rho&\mathcal{C}\\ \rho&1&0\\ \mathcal{C}&0&1\end{array}\right),$ (D.2)
@@
$1-\rho^{2}-\mathcal{C}^{2}\geq 0,$ (D.3)
which, when plugged in (D.1), gives (3.4).
@@
Appendix E Experimental details
@@
$\Sigma_{xy}=\Sigma_{yx}=(\Sigma_{yy}-\beta I)^{1/2},$ (E.1)

### v2

@@ B.1 Proofs on $S^{\Sigma}_{x}$
$=x^{\top}\left(\Sigma_{xx}^{-1}+\Sigma_{xx}^{-1}\Sigma_{xy}{S^{\Sigma}_{x}}^{-1}\Sigma_{yx}\Sigma_{xx}^{-1}\right)x+y^{\top}{S^{\Sigma}_{x}}^{-1}y+x^{\top}\left(-\Sigma_{xx}^{-1}\Sigma_{xy}{S^{\Sigma}_{x}}^{-1}\right)y+y^{\top}\left(-{S^{\Sigma}_{x}}^{-1}\Sigma_{yx}\Sigma_{xx}^{-1}\right)x$
@@ B.2 Remarks on Assumption 1
Let Assumptions 4 and 6 hold, and let $n=\Theta(d)$. Let $\hat{\theta}_{\textup{LR}}(\lambda)$ be defined as in (4.4), and let $\mathcal{C}(\hat{\theta}_{\textup{LR}}(\lambda))$ be the amount of spurious correlations learned by the model $f_{\textup{LR}}\left(\hat{\theta}_{\textup{LR}}(\lambda),\cdot\right)$ as defined in (3.3). Then, for any $\lambda>0$, we have that, for every $t\in(0,1/2)$,
@@
Let Assumptions 4 and 6 hold, and let $n=\Theta(d)$. Let $\hat{\theta}_{\textup{LR}}(\lambda)$ be defined as in (4.4), and let $\mathcal{L}(\hat{\theta}_{\textup{LR}}(\lambda))$ be the in-distribution test loss of the model $f_{\textup{LR}}(\hat{\theta}_{\textup{LR}}(\lambda),\cdot)$ as defined in (4.1). Then, , for any $\lambda>0$, we have that, for every $t\in(0,1/2)$,
@@ Appendix C Proofs for Random Features
$\begin{split}\lambda_{\rm min}\left(\Phi\Phi^{\top}\right)&\geq p\,\lambda_{\rm min}\left(\mu_{1}^{2}\frac{ZZ^{\top}}{2d}+{\tilde{\mu}}^{2}I\right)-\left\lVert\mathbb{E}_{V}\left[\Phi\Phi^{\top}\right]-p\left(\mu_{1}^{2}\frac{ZZ^{\top}}{2d}+{\tilde{\mu}}^{2}I\right)\right\rVert_{\textup{op}}-\left\lVert\Phi\Phi^{\top}-\mathbb{E}_{V}\left[\Phi\Phi^{\top}\right]\right\rVert_{\textup{op}}\\ &\geq p{\tilde{\mu}}^{2}-\mathcal{O}\left(\frac{p\log^{3}d}{\sqrt{d}}\right)-\mathcal{O}\left(\sqrt{pd}\right)=\Omega(p),\end{split}$ (C.30)
@@
Let $z\in\mathbb{R}^{2d}$ be sampled from a distribution respecting Assumption 4, not necessarily with the same covariance as $\mathcal{P}_{XY}$, independent from everything else, and let $f_{\textup{RF}}(\hat{\theta}_{\textup{RF}}(\lambda),z)$ be the RF model defined in (6.1) with $\hat{\theta}_{\textup{RF}}(\lambda)$ defined in (6.3), with $\lambda\geq 0$. Then, we have that
$\left|f_{\textup{RF}}(\hat{\theta}_{\textup{RF}}(\lambda),z)-\mu_{1}^{2}p\frac{z^{\top}Z^{\top}}{2d}\left(\Phi\Phi^{\top}+n\lambda I\right)^{-1}G\right|=\mathcal{O}\left(\frac{d^{1/4}\log d}{p^{1/4}}+\frac{\log^{3/2}d}{d^{1/8}}\right)=o(1),$ (C.58)
@@
$f_{\textup{RF}}(\hat{\theta}_{\textup{RF}}(\lambda),z)$ $=\left(\mu_{1}Vz+\tilde{\phi}\left(Vz\right)\right)^{\top}\left(\mu_{1}VZ^{\top}+\tilde{\Phi}^{\top}\right)\hat{G}$ (C.59)
@@
$\left|f_{\textup{RF}}(\hat{\theta}_{\textup{RF}}(\lambda),z)-\mu_{1}^{2}p\frac{z^{\top}Z^{\top}}{2d}\hat{G}\right|=\mathcal{O}\left(\frac{d^{1/4}\log d}{p^{1/4}}+\frac{\log^{3/2}d}{d^{1/8}}\right),$ (C.68)
@@
$=f_{\textup{LR}}(\hat{\theta}_{\textup{LR}}(\tilde{\lambda}),z),$
@@
$\left|\mu_{1}^{2}p\frac{z^{\top}Z^{\top}}{2d}\left(\Phi\Phi^{\top}+n\lambda I\right)^{-1}G-f_{\textup{RF}}(\hat{\theta}_{\textup{RF}}(\lambda),z)\right|=\mathcal{O}\left(\frac{d^{1/4}\log d}{p^{1/4}}+\frac{\log^{3/2}d}{d^{1/8}}\right),$ (C.75)
@@
Appendix D Features Composed as $z=x+y$
In this work, we consider the data to be composed as $z=[x^{\top},y^{\top}]^{\top}$. However, some of the high level results we obtain can be recovered for settings where the features are composed differently. As an example, let us consider the model
$z=x+y,\qquad g=x^{\top}\theta^{*}+\varepsilon.$ (D.1)
Then, according to (3.3), we have
$\mathcal{C}(\hat{\theta})=\mathrm{Cov}\left((\tilde{x}+y)^{\top}\hat{\theta},x^{\top}\theta^{*}\right)=\hat{\theta}^{\top}\Sigma_{yx}\theta^{*},$ (D.2)
which provides a quantity that could be studied again via the analysis in [19]. In this new setup, we remark that the full data covariance takes the form
$\Sigma_{zz}=\Sigma_{xx}+\Sigma_{yy}+\Sigma_{xy}+\Sigma_{yx}.$ (D.3)
In a nutshell, we expect the analysis for this setting to provide a qualitative behavior similar to that unveiled by the case $z=[x^{\top},y^{\top}]^{\top}$. In fact, the experiments on Color-MNIST (which does not strictly follow the model $z=[x^{\top},y^{\top}]^{\top}$, as the color overlaps with the core feature pattern as in the model $z=x+y$) suggest that our conclusions hold beyond the setting of “orthogonal features”. We further remark that in the setting $z=[x^{\top},y^{\top}]^{\top}$ the optimal solution $\hat{\theta}=\theta^{*}$ gives $\mathcal{C}=0$, while this is not necessarily the case in the setting $z=x+y$.
Appendix E Proof of (3.4): connection with out-of-distribution loss
@@ Appendix D Features Composed as $z=x+y$
$\mathbb{E}_{\tilde{x},y}\left[\left(f(\hat{\theta},[\tilde{x}^{\top},y])-f^{*}_{x}(\tilde{x})\right)^{2}\right]$ $=2-2\mathbb{E}_{\tilde{x},y}\left[f(\hat{\theta},[\tilde{x}^{\top},y])f^{*}_{x}(\tilde{x})\right]$ (E.1)
@@
$S=\left(\begin{array}[]{@{}c c c@{}}1&\rho&\mathcal{C}\\ \rho&1&0\\ \mathcal{C}&0&1\end{array}\right),$ (E.2)
@@
$1-\rho^{2}-\mathcal{C}^{2}\geq 0,$ (E.3)
which, when plugged in (E.1), gives (3.4).
@@
Appendix F Experimental details
@@
$\Sigma_{xy}=\Sigma_{yx}=(\Sigma_{yy}-\beta I)^{1/2},$ (F.1)
@@
F.1 Additional Experiments
Figure 7: Test loss $\mathcal{L}(\hat{\theta}_{\textup{NN}/\textup{LR}}(\lambda))$ (black) and spurious correlations $\mathcal{C}(\hat{\theta}_{\textup{NN}/\textup{LR}}(\lambda)$ (red) as a function of $\lambda$. First and second panel: 2-layer fully connected ReLU network, trained on the multi-class color(C)-MNIST, for two different values of $\alpha$. Third and fourth panel: Same setup as in Figure 2, for the binary C-MNIST dataset, with multiple values of $\alpha$.
Figure 8: Test loss $\mathcal{L}(\hat{\theta}_{\textup{NN}/\textup{RF}}(\lambda))$ (black) and spurious correlations $\mathcal{C}(\hat{\theta}_{\textup{NN}/\textup{RF}}(\lambda)$ (red) as a function of $\lambda$. Left: 2-layer fully connected ReLU network, trained on the Corrupted CIFAR-10 dataset (boats and trucks, with added textures “brightness” and “glass blur”). Right: RF model with $\tanh$.
In the left two panels of Figure 7 we consider Color-MNIST including all 10 digits. We train a 2-layer network on all classes with one-hot encoding and MSE loss. Odd (even) digits are red (blue) with probability $(1+\alpha)/2$, and blue (red) with probability $(1-\alpha)/2$. To compute $\mathcal{C}$, at test time we consider the parity of the logit with the highest value, with respect to the color of the image. The first and the second panel correspond to two values of $\alpha$-s and follow a similar profile as Figure 5 (left), showing the same qualitative behavior of $\mathcal{L}$ and $\mathcal{C}$ with respect to $\lambda$ for the full Color-MNIST dataset (i.e., in the multi-class setting). In the third and fourth panel of Figure 7 we repeat the experiment in Figure 2 (right) for multiple values of $\alpha$, reporting $\mathcal{L}$ and $\mathcal{C}$ with respect to $\lambda$ for linear regression. The curves behave as expected: for any value of $\lambda$, as $\alpha$ decreases, $\mathcal{C}$ decreases. Furthermore, the (in-distribution) test loss decreases as $\alpha$ increases, in agreement with our discussion at the end of Section 5.
In Figure 8 we train a 2-layer network and a random feature model on the classes “trucks” and “boats” from the Corrupted CIFAR-10 dataset (used in the context of spurious correlations in [30]), enforcing a correlation in the training set with respectively the textures “brightness” and “glass blur”, with a correlation $\alpha=0.95$. In both panels, we see a mild increase in $\mathcal{C}$ as $\lambda$ initially increases, until the later decrease predicted by Proposition 5.1. The profiles are also qualitatively similar to the ones of Figure 5 (left).